OpenStudy (anonymous):

How to integrate ; 2/3 ∫cos(x)dx+2/3 ∫cos(5x)dx

OpenStudy (anonymous):

Is this integration direct . No need to use U-substitution rite...

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

We can use U-substitution for the second integrand, but after enough experience you'll begin to see that \[\int\limits \cos(5x)dx = {1 \over 5} \sin(5x)\]

OpenStudy (anonymous):

2/3 ∫cos(x)dx+2/3 ∫cos(5x)dx=2/3 sinx + 2/15 sin5x +c

OpenStudy (anonymous):

Thanks guys....